OpenStudy (anonymous):

1. Solve: (4x - 5)(3x + 2) =0 2. Factor: x2 + 11x + 10

OpenStudy (mayliz):

1. 4x-5=0 x=5/4 or 3x+2=0 x=-2/3

OpenStudy (anonymous):

do you kno the second one?

OpenStudy (mayliz):

x2+11x+10=0?

OpenStudy (anonymous):

ya

OpenStudy (mayliz):

d=121+40=161